OpenStudy (anonymous):

@hashsam1 has 3 yellow, 8 green, 5 red, and 14 blue ankle bracelets in a bag. If hashsam1 randomly selects an ankle bracelet out of the bag, what is the probability it will be green or blue?

OpenStudy (anonymous):

hahahahahahaha :DDDDDD

OpenStudy (anonymous):

i won on ur last question rite. say yes.

OpenStudy (anonymous):

woohooo

OpenStudy (anonymous):

A.6/9 B.11/15 C. 3/14 D. 3/4

OpenStudy (anonymous):

when ever there is or in the probability then you need to add the two different probabilities if its green= green/total = 8/30 = 4/15 if its blue = blue/total = 14/30=7/15 add both 7/15 + 4/15= 11/15 B

OpenStudy (anonymous):

this dudes good^

OpenStudy (anonymous):

:D if you are referring to me Ty

OpenStudy (anonymous):

welcome! (: